The food we eat is degraded,or broken down, in our bodies to provide energy for growth amd function. A general overall question for this very complex process represents the degradation of glucose ( C6 H12 O6) to carbon dioxide (CO2) and water ( H2O)

C6 H12 06 arrow 6CO2 + 6H2O

If 856 g of C6 H12 O6 is consumed by a person over a certain period, what is the mass of CO2 produced?
1
Expert's answer
2018-04-25T08:38:36-0400
m(C6H12O6) = 856 g

M(C6H12O6) = 12*6+1*12+16*6 = 180 g/mol

n(C6H12O6) = m/M = 856 g / 180 g/mol = 4.756 mol

n(CO2) = 6*4.756 = 28.536 mol

M(CO2) = 12+16*2 = 44 g/mol

m(CO2) = n*M = 28.536mol*44g/mol = 1255.584 g